# ON THE AVERAGE ULTRAVIOLET EMISSION LINE SPECTRA OF HIGH-REDSHIFT GALAXIES: HOT AND COLD, CARBON-POOR, NITROGEN-MODEST, AND OOZING IONIZING PHOTONS

### v1

@@ 2.3 Emission Line Measurements
1487 | N iv] | blended doublet |
1550 | C iv | blended resonance doublet |
1640 | He ii | – |
1663 | O iii] + | blended pair, marginally |
@@
2880 | Mg ii | blended resonance doublet |
3727 | [O ii] | blended doublet |
3869 | [Ne iii] | – |
3889 | He i | marginally resolved |
@@
3970 | [Ne iii] +H | blended pair, unresolved |
4102 | H | – |
4341 | H | – |
4363 | [O iii] | marginally resolved |
@@
4687 | He ii 4687 | – |
4862 | H | – |
4960 | [O iii] | – |
5008 | [O iii] | 4960 doublet companion |
5877 | He i | – |
6302 | [O i] | – |
6564 | H + | blended complex, |
@@
For each set of stacked spectra, we measure the lumiosities (in the continuum-subtracted spectra) and EWs (in the normalized spectra) for a large number of UV and optical lines. Some lines, particularly a number of UV doublets, are blended at the resolution of NIRSpec’s PRISM, and cannot be measured independently. We list the lines we fit in Table 2, providing comments where appropriate.
@@
In the UV, we are unable to deblend doublets of C iv, O iii], C iii], Mg ii, or the N iii] multiplet (1747, 1749, 1750, 1752, and 1754Å). Importantly, however, we are able to deblend the He ii line from the O iii] doublet – these lines are separated by restframe Å, which corresponds to marginally more than one resolution element. In the optical, we are unable to deblend the doublets of [O ii] and [S ii], as well as blended pairs of [N ii]+H, and [Ne iii]+H. We fit all these blends with a single emission line. The [O iii]4960,5008 lines are both resolved from each other and clearly separated from H, as well as [Ne iii] from He i and H from [O iii]. See Table 2 for a complete list.
@@
We fit individual emission lines using single Gaussian profiles and applying the lmfit minimization algorithm (Newville et al., 2016) to both continuum-subtracted (to get absolute line luminosities) and continuum-normalized -stacked spectra (for EWs). Free parameters are the line amplitude (in either luminosity or normalized units), the central rest-frame wavelength (in Å, see priors in Table 2) and the width of each line (in Å). We also record the luminosity and EW of each line by numerical integration, using a 5,000 km s-1custom window in the observed frame, approximately.
@@
The amplitude is assumed to be positive () for all lines except for C iv, where we also allow for negative values () in order to account for the stellar P-Cygni+ISM components of this line (in absorption, rather than nebular dominated). Emission lines within blended doublets or pairs are fitted simultaneously, and the wavelength separation between them is fixed to the value given by the rest-frame wavelength of the constituents. Close doublets and multiplets are fitted with single component Gaussians at the average wavelength – see Table 2 for details.
@@
Errors on the line parameters were derived using a Monte-Carlo approach, perturbing every pixel-luminosity by a Gaussian distribution whose mean is zero and standard deviation is the error coming from the bootstrap simulations, then re-fitting each line over iterations per stacked spectrum. The resulting line luminosities and EWs are presented in Sect. 3.1, and discussed in the context of this paper.
3 Results I: The Average Spectrum of Galaxies
@@
In Figure 2 we show the average spectra of the galaxies in our sample, when binned in five groups sorted along a vector of the [O iii] 5008 equivalent width (). probes, on average, the youth of the star-forming episode by tracing the ionizing photon production rate, modulated by the ionization parameter and the nebular oxygen abundance. The highest values are found among galaxies hosting very young, intense starbursts with weak stellar continua, and . [O iii] is often the brightest line in the optical spectrum of starbursts, and visible with NIRSpec/PRISM out to . These bins have median values of ranging from 120Å for the lowest EW bin and 2200Å for the highest.
stack # | C IV/1550 | He II/1640 | [O III]/1666 | C III/1909 | [C II]/2326 | MgII/2800 |
1 | ||||||
2 | ||||||
3 | ||||||
4 | ||||||
5 |
@@
1 | ||||||||
2 | ||||||||
3 | ||||||||
4 | ||||||||
5 |
stack # | C IV/1550 | He II/1640 | [O III]/1666 | C III/1909 | [C II]/2326 | MgII/2800 |
1 | ||||||
2 | ||||||
3 | ||||||
4 | ||||||
5 |
@@ 3.1 Evolution of the Average UV Spectrum with
– Emission lines of C iv, the blend of He ii and O iii], and C iii] are all particularly clear, and seen in emission in the stacks of highest (e.g. yellow curve in the lower left panel of Figure 2). The higher excitation energies of these UV lines compared to the optical lines makes them detectable only at the high temperatures associated with low gas-phase metallicities (e.g. Mingozzi et al., 2022). In contrast, none of these features are seen in emission for the lowest bins (dark purple curve in the same figure). Together with optical emission lines, these features can be used to estimate the electron temperatures and abundances of oxygen and carbon in the interstellar medium (see Section 6.3).
@@
– There are detections of both the N iv] and N iii] multiplet in the highest stacks. These can provide estimates of the nitrogen abundance which in turn can point towards the presence of exotic stellar populations. We discuss this in more detail in dedicated stacks in Section 6.3.
@@
We compare the stacked NIRSpec spectra with spectra obtained in the low- universe, for which we adopt the COS Legacy Archive Spectroscopic SurveY (CLASSY; Berg et al. 2022; James et al. 2022). The CLASSY sample comprises 45 vigorously star forming galaxies at , for which deep and almost-homogeneous spectroscopy has been obtained with HST/COS in all of the medium resolution gratings. The primary selection criteria of CLASSY extracted a sample comparable in various characteristics (SFR, stellar mass, size, oxygen abundance, etc.) to those identified at redshifts above 6 (see Berg et al., 2022, for details). We treat these spectra similarly to those of the NIRSpec sample, first removing the continuum through subtraction and normalization, then smoothing them to the resolving power of NIRSpec appropriate for the wavelength. We then stack all the CLASSY sources with Å, which we plot as the black lines in the lower left panels of Figure 2. We note that CLASSY’s COS spectra have been observed in the G185M grating to measure the C iii]1909 line, but this grating has a very narrow spectral bandwidth which leads to a gap just bluewards of this line before the G160M grating. We have no UV spectroscopy at redder wavelengths, so the [C ii] multiplet and Mg ii emission are unknown in CLASSY for the moment.
@@ 3.2 Quantitative Trends and Comparisons
We measure the strength of the UV emission lines, as described in Section 2.3. We show the evolution of some characteristic EWs in Figure 3, and present the fluxes and equivalent widths in Tables 3, 4, 5, and 6. We assess their evolution on various diagnostic diagrams, and contrast the high- average spectra with other individual targets, in Figure 4. This Figure shows diagrams of (a.) the C iii]-EW vs. the C iii/He ii] ratio (left-most), (b.) the C iii]/He ii vs. the O iii]/He ii ratio (center-left), (c.) the C iv/C iii] ratio vs. the (C iii+C iv)/He ii ratio (center-right), and (d.) C iv/C iii] ratio vs. O iii]/He ii (right-most). When galaxies are ordered by starburst intensity (increasing ), it is clearly visible how the C iii] EW also increases, rising from Å to 15Å for the most intense starburst (e.g. central panel). This evolution in EW is unsurprising, since these lines are, to first order, connected to the rate of ionizing photon production at energies above and eV, respectively, and naturally we expect their EWs to broadly correlate. However, the two right-most plots show that the C iv/C iii] ratio also increases along the same sequence of , implying there is also an increase in the hardness of the ionizing spectrum (probed here by ionization edges at 45 and 24 eV).
Light green points show the position of galaxies from the CLASSY sample (Berg et al., 2022; James et al., 2022), including only galaxies for which all these flux measurements can be made individually (i.e. showing no limits). This paints a different picture to the qualitative discussion based on the stacked CLASSY spectra (Section 3.1): the C iv/C iii] ratios of the stacked high- spectra reach values of , while the median of green points is higher, and reaches values of almost 1. This happens despite the fact that both lines must be detected to participate in the diagram, the real dispersion in the low- sample remains very large compared to the averages of the sample.
AGN, star-forming, and shocked gas regions are overplotted from models of Feltre et al. (2016), Allen et al. (2008), and Gutkin et al. (2016) in the line ratio diagrams, as well as from Nakajima et al. (2018). It is clear that our stacks, as well as all the CLASSY galaxies, are well confined to the star-formation dominated regions in the line ratios and do not require additional sources of ionization. However, the two highest stacks, as well as a handful of CLASSY galaxies do encroach in the AGN region when is used on one of the axes. See Section 6 for more discussion on this.
@@ 4 Results II: Electron Temperatures and Metal Abundances
The spectra show various nebular emission lines from the O++ ion, including the strong lines at 4960 and 5008 Å, the auroral line at 4363 Å, and a blend of the intercombination lines at 1661 and 1666 Å in the ultraviolet. We also see a significant change in the amplitude of the Balmer Jump (BJ) near 3650 Å (Figure 2, lower panels), which we highlight in Figure 5. The ratios of [O iii]1666/5008 and [O iii]4363/5008, as well as the amplitude of the Balmer Jump, all provide estimates of the characteristic kinetic energy of the electrons in the ionized gas, and therefore the electron temperature, . Because most of the cooling in nebular gas occurs through collisions with ionized metals (primarily O+ and O++), this temperature is fundamentally influenced by the metal abundance. We now explore each of these topics and measurements in turn.
@@ 4.1.1 Balmer Jump
One of the most noteworthy features of the optical spectrum in Figure 2 is the presence of a Balmer jump, for which we show a zoom-in in Figure 5. In the lower stacks (purple, blue) this is positive, and arises because A and B type stars dominate the light at Å: these stars have suitable temperatures to produce a significant population of neutral hydrogen with electrons excited to the level. Photoionization in the stellar atmosphere out of this state produces the stellar Balmer break. In contrast, the higher stacks (orange, yellow) show an inverted break: this arises because of free electrons in the H ii regions, which can recombine into the or (or other) states, producing the Balmer and Paschen (or other) continua, respectively. The amplitude of this break depends on the ratio of the densities of electrons with and eV, and hence is temperature dependent: the amplitude of the nebular Balmer break decreases with increasing .
Our spectra contain both stellar and nebular emission, and we therefore cannot simply measure the amplitude of the nebular Balmer jump to derive , because of the unknown contribution of the stellar jump. We therefore build a spectral model to infer , accounting for both stellar continuum, and nebular radiation from the recombination lines, free-bound, free-free, and 2-photon continua of H and He. For starlight we take template spectra from Starburst99 (Leitherer et al., 1999; Vázquez & Leitherer, 2005), assuming stellar evolutionary tracks from the Geneva group, a Salpeter IMF between 0.1 and 100 M⊙ and all available metallicities. We fit an age parameter assuming a constant star formation rate (SFR), using the Reddy et al. (2015, 2016) dust attenuation laws, and an arbitrary normalization parameter (as the spectra are normalized). The nebular gas is generated using the code of Schirmer (2016), which we run on a grid spaced by 1000 K and then interpolate using spline functions. The nebular reddening is treated across the full wavelength range, again using the Reddy et al. (2015, 2016) law; this is the same law as used for the stellar light although stellar and nebular reddenings are allowed to be independent. The model conserves the number of ionizing photons and normalizes the spectrum to the total (maximum) photoionization, but introduces another scale factor to account for a fraction of ‘lost’ LyC photons, which either escape the galaxy or are absorbed by dust grains.
@@
An example of this method is shown in the left panel of Figure 6. Here we show the data from the stack of the highest galaxies in gray, and the best-fitting pure stellar model is shown in black. This stellar fit well captures the very blue end of the UV ( Å) and the red end of the optical ( Å), but does not match the data between in the vicinity of the Balmer break. The cyan lines, however, represent a combined fit of stellar continuum and nebular gas, constrained by the Balmer lines and total continuum level: this almost perfectly reproduces the complete spectrum, including the region around the Balmer jump. The right hand panel of the same Figure shows compared to the age of the stellar population from the same fit. Temperatures are found to range from 12,000–13,000 K, which are found preferentially in galaxies with log(Age[Myr]) , down to temperatures as low as 8,000 K, which are found in galaxies with log(Age[Myr]) –8 Myr. These temperatures are reported in Table 7.

### v2

@@ 2.3 Emission Line Measurements
| 1487 | N iv] | blended doublet |
| 1550 | C iv | blended resonance doublet |
| 1640 | He ii | – |
| 1663 | O iii] + | blended pair, marginally |
@@
| 2880 | Mg ii | blended resonance doublet |
| 3727 | [O ii] | blended doublet |
| 3869 | [Ne iii] | – |
| 3889 | He i | marginally resolved |
@@
| 3970 | [Ne iii] +H | blended pair, unresolved |
| 4102 | H | – |
| 4341 | H | – |
| 4363 | [O iii] | marginally resolved |
@@
| 4687 | He ii 4687 | – |
| 4862 | H | – |
| 4960 | [O iii] | – |
| 5008 | [O iii] | 4960 doublet companion |
| 5877 | He i | – |
| 6302 | [O i] | – |
| 6564 | H + | blended complex, |
@@
For each set of stacked spectra, we measure the luminosities (in the continuum-subtracted spectra) and EWs (in the normalized spectra) for a large number of UV and optical lines. Some lines, particularly a number of UV doublets, are blended at the resolution of NIRSpec’s PRISM, and cannot be measured independently. We list the lines we fit in Table 2, providing comments where appropriate.
@@
In the UV, we are unable to deblend doublets of C iv, O iii], C iii], Mg ii, or the N iii] multiplet (1747, 1749, 1750, 1752, and 1754Å). Importantly, however, we are able to deblend the He ii line from the O iii] doublet – these lines are separated by restframe Å, which corresponds to marginally more than one resolution element. This can be seen by contrasting the He ii+[O iii] complex with the C iv doublet in the lower left panel of Figure 2: the difference in spectral resolution is small across this wavelength range, but the He ii+[O iii] features are clearly broader than C iv, by a factor of around 2. We impose a fixed wavelength separation between He ii and the O iii] blend (see Table 2), constrain the FWHM to be equal, and simultaneously fit for the amplitude of two Gaussian components. In the optical, we are unable to deblend the doublets of [O ii] and [S ii], as well as blends of [N ii]+H, and [Ne iii]+H. We fit all these blends with a single emission line. The [O iii]4960,5008 lines are both resolved from each other and clearly separated from H, as well as [Ne iii] from He i and H from [O iii]. See Table 2 for a complete list.
@@
We fit individual emission lines using single Gaussian profiles and applying the lmfit minimization algorithm (Newville et al., 2016) to both continuum-subtracted (to get absolute line luminosities) and continuum-normalized -stacked spectra (for EWs). Free parameters are the line amplitude, the central rest-frame wavelength (in Å, see priors in Table 2) and the width of each line (in Å). We also record the luminosity and EW of each line by numerical integration, using a 5,000 km s-1 window.
@@
The amplitude is constrained to be positive () for all lines except for C iv and Mg ii, for which we also allow for negative values () in order to account for stellar and ISM absorption components. Close-by emission lines within blended doublets or pairs are fitted simultaneously, and the wavelength separation between them is fixed to the value given by the rest-frame wavelength of the constituents. Close doublets and multiplets are fitted with single component Gaussians at the average wavelength – see Table 2 for details.
@@
Errors on the line parameters were derived using a Monte-Carlo approach, perturbing the value of every spectral pixel by a Gaussian distribution whose mean is zero and standard deviation is the error coming from the bootstrap simulations, then re-fitting each line over iterations per stacked spectrum. The resulting line luminosities and EWs are presented in Sect. 3.1, and discussed in the context of this paper.
3 Results I: The Average Spectrum of Galaxies at Redshift Beyond 4
@@
In Figure 2 we show the average spectra of the galaxies in our sample, when binned in five groups sorted along a vector of . probes, on average, the youth of the star-forming episode by tracing the ionizing photon production rate, modulated by the ionization parameter and the nebular oxygen abundance. The highest values are found among galaxies hosting very young, intense starbursts with weak stellar continua, and (e.g. Izotov et al. 2023 and the sample of Hayes et al. 2023, at least at low-). [O iii] is often the brightest line in the optical spectrum of starbursts, and visible with NIRSpec/PRISM out to . These bins have median values of ranging from 120Å for the lowest EW bin and 2200Å for the highest (See Table 1).
| stack # | Civ/1550 | Heii/1640 | [Oiii]/1666 | Ciii]/1909 | [Cii]/2326 | Mgii/2800 |
| 1 | ||||||
| 2 | ||||||
| 3 | ||||||
| 4 | ||||||
| 5 |
@@
| 1 | ||||||||
| 2 | ||||||||
| 3 | ||||||||
| 4 | ||||||||
| 5 |
| stack # | Civ/1550 | Heii/1640 | [Oiii]/1666 | Ciii]/1909 | [Cii]/2326 | Mgii/2800 |
| 1 | ||||||
| 2 | ||||||
| 3 | ||||||
| 4 | ||||||
| 5 |
@@ 3.1 Evolution of the Average UV Spectrum with
– Emission lines of C iv, the blend of He ii and O iii], and C iii] are all particularly clear, and seen in emission in the stacks of highest (e.g. yellow curve in the lower left panel of Figure 2). The higher excitation energies of these UV lines compared to most optical lines observed at high- makes them detectable only at the high temperatures associated with low gas-phase metallicities (e.g. Mingozzi et al., 2022). In contrast, none of these features are seen in emission for the lowest bins (dark purple curve in the same figure). Together with optical emission lines, these features can be used to estimate the electron temperatures and abundances of oxygen and carbon in the interstellar medium (see Section 6.3).
@@
– There are detections of both the N iv] and N iii] multiplet in the highest stacks. These can provide estimates of the nitrogen abundance which in turn can point towards the presence of exotic stellar populations. We discuss this in more detail in Section 6.3.
@@
We also compare the stacked NIRSpec spectra with spectra obtained in the low- universe, for which we adopt the COS Legacy Archive Spectroscopic SurveY (CLASSY; Berg et al. 2022; James et al. 2022). The CLASSY sample comprises 45 vigorously star forming galaxies at , for which deep and almost-homogeneous spectroscopy has been obtained with HST/COS in all of the medium resolution gratings. The primary selection criteria of CLASSY extracted a sample comparable in various characteristics (SFR, stellar mass, size, oxygen abundance, etc.) to those identified at redshifts above 6 (see Berg et al., 2022, for details). We treat these spectra similarly to those of the NIRSpec sample, first removing the continuum through subtraction and normalization, then smoothing them to the resolving power of NIRSpec appropriate for the wavelength. We then stack all the CLASSY sources with Å, which we plot as the black lines in the lower left panels of Figure 2. We note that CLASSY’s COS spectra have been observed in the G185M grating to measure the C iii]1909 line, but this grating has a very narrow spectral bandwidth which leads to a gap just bluewards of this line before the G160M grating. We have no UV spectroscopy at redder wavelengths, so the [C ii] multiplet and Mg ii emission are unknown in CLASSY for the moment.
@@ 3.2 Quantitative Trends and Comparisons
We measure the strength of the UV emission lines, as described in Section 2.3. We show the evolution of some characteristic EWs in Figure 3, and present the fluxes and equivalent widths in Tables 3, 4, 5, and 6. Several results are based upon line ratios, of which the major ratios used in this paper are listed in Table 7. We first correct all emission lines for differential dust obscuration using the Reddy et al. (2015, 2016) attenuation law, and dust reddening estimated from the H/H line ratio assuming an intrinsic value of 2.86, which is appropriate for K gas.
We assess their evolution on various diagnostic diagrams, and contrast the high- average spectra with other individual targets, in Figure 4. This Figure shows diagrams of (a.) the C iii]-EW vs. the C iii]/He ii ratio (left-most), (b.) the C iii]/He ii vs. the O iii]/He ii ratio (center-left), (c.) the C iv/C iii] ratio vs. the (C iii]+C iv)/He ii ratio (center-right), and (d.) C iv/C iii] ratio vs. O iii]/He ii (right-most). When galaxies are ordered by starburst intensity (increasing ), it is clearly visible how the C iii] EW also increases, rising from Å to 15Å for the most intense starburst (e.g. central panel). This evolution in EW is unsurprising, since these lines are, to first order, connected to the rate of ionizing photon production at energies above and eV, respectively, and naturally we expect their EWs to broadly correlate. However, the two right-most plots show that the C iv/C iii] ratio also increases along the same sequence of , implying there is also an increase in the hardness of the ionizing spectrum (probed here by ionization edges at 45 and 24 eV).
Light green points show the position of galaxies from the CLASSY sample (Berg et al., 2022; James et al., 2022), including only galaxies for which all these flux measurements can be made individually (i.e. showing no limits). This paints a different picture to the qualitative discussion based on the stacked CLASSY spectra (Section 3.1): the C iv/C iii] ratios of the stacked high- spectra reach values of , while the median of green points is higher, and reaches values of almost 1. This happens despite the fact that both lines must be detected to appear on the diagram, and the real dispersion in the low- sample remains very large compared to the averages of the sample.
AGN, star-forming, and shocked gas regions are overplotted from models of Feltre et al. (2016), Allen et al. (2008), and Gutkin et al. (2016) in the line ratio diagrams, as well as demarcated from Nakajima et al. (2018). It is clear that our stacks, as well as all the CLASSY galaxies, are well confined to the star-formation dominated regions in the line ratios and do not require additional sources of ionization. However, the two highest stacks, as well as a handful of CLASSY galaxies, do encroach into the AGN region when is used on one of the axes. See Section 6 for more discussion on this.
@@ 4 Results II: Electron Temperatures and Metal Abundances
The spectra show various nebular emission lines from the O++ ion, including the strong lines at 4960 and 5008 Å, the auroral line at 4363 Å, and a blend of the intercombination lines at 1661 and 1666 Å in the ultraviolet. We also see a significant change in the amplitude of the Balmer Jump (BJ) near 3650 Å (Figure 2, lower panels), which we highlight in Figure 5. The ratios of [O iii]5008/1666 and [O iii]5008/4363, as well as the amplitude of the Balmer Jump, all provide estimates of the characteristic kinetic energy of the electrons in the ionized gas, and therefore the electron temperature, . Because most of the cooling in nebular gas occurs through collisions with ionized metals (primarily O+ and O++), this temperature is fundamentally influenced by the metal abundance. We now explore each of these topics and measurements in turn.
@@ 4.1.1 Balmer Jump
. Strong emission lines are labeled. The color coding is the same as used in Figure 2.
One of the most noteworthy features of the optical spectra of Figure 2 is the presence of a Balmer jump, which is clearly visible in the two stacks with the highest , and for which we show a zoom-in in Figure 5. In the lower stacks (purple, blue) this break is positive, and arises because A and B type stars contribute significantly to the light at Å: these stars have suitable temperatures to produce a significant population of neutral hydrogen with electrons excited to the level. Photoionization in the stellar atmosphere out of this state produces the stellar Balmer break. In contrast, the higher stacks (orange, yellow) show an inverted break: this arises because of free electrons in the H ii regions, which can recombine into the or (or other) states, producing the Balmer and Paschen (or other) continua, respectively. The amplitude of this break depends on the ratio of the densities of electrons with and eV, and hence the gas temperature: the amplitude of the nebular Balmer jump decreases with increasing .
Our spectra contain both stellar and nebular emission, and we therefore cannot simply measure the amplitude of the nebular Balmer jump to derive , because of the unknown contribution of the stellar break. We therefore build a spectral model to infer , accounting for both stellar continuum, and nebular radiation from the recombination lines, free-bound, free-free, and 2-photon continua of H and He. For starlight we take template spectra from Starburst99 (Leitherer et al., 1999; Vázquez & Leitherer, 2005), assuming stellar evolutionary tracks from the Geneva group, a Salpeter IMF between 0.1 and 100 M⊙ and all available metallicities. We fit an age parameter assuming a constant star formation rate (SFR), using the Reddy et al. (2015, 2016) dust attenuation laws, and an arbitrary normalization parameter (as the spectra are normalized). The nebular gas is generated using the code of Schirmer (2016), which we run on a grid spaced by 1000 K and then interpolate using spline functions. The nebular reddening is treated across the full wavelength range, again using the Reddy et al. (2015, 2016) law; this is the same law as used for the stellar light although stellar and nebular reddenings are allowed to be independent. The model conserves the number of ionizing photons and normalizes the spectrum to the total (maximum) photoionization, but introduces another scale factor to account for a fraction of ‘lost’ LyC photons, which either escape the galaxy or are absorbed by dust grains.
@@
An example of this method is shown in the left panel of Figure 6. Here we show the data from the stack of the highest galaxies in gray, and the best-fitting pure stellar model is shown in black. This stellar fit well captures the very blue end of the UV ( Å) and the red end of the optical ( Å), but does not match the data in the vicinity of the Balmer break. The cyan lines, however, represent a combined fit of stellar continuum and nebular gas, constrained by the Balmer lines and total continuum level: this almost perfectly reproduces the complete spectrum, including the region around the Balmer jump. The right hand panel of the same Figure shows compared to the age of the stellar population from the same fit. Temperatures are found to range from ,000 K, which are found preferentially in galaxies with ages Myr, down to temperatures as low as 8,000 K, which are found in galaxies with ages of to yr. These temperatures are reported in Table 8.